OpenStudy (anonymous):

Assume that y varies directly as x. If y = -8 when x = -3, find x when y = 6. Leave your answer in fractional form. help please!!

OpenStudy (jordan):

-2/3

OpenStudy (anonymous):

X = ky -3 = -8k K = 3/8 x = ky X = 3/8 * 6 X = 2.25

OpenStudy (anonymous):

Or 9/4

OpenStudy (jordan):

X = ky -3 = -8k K = 3/8 x = ky X = 3/8 * 6 X = 2.2555

OpenStudy (anonymous):

thanks (: